Exercise 11. For the joint distribution given in Table 3, find the variance of W = X + Y directly, and also by using Theorem 5.6.


Theorem 5.6. Given random variables X and Y, and constants a and b, VarſaX + bY] = aʼVar[X] + b2V ar[Y] + 2ab cov(X,Y)
-3: Joint Distribution of X, Y Prob Y = 1 Y = 2 X=1 0.04 0.36 X = 2 0.16 0.44
Answer #1

w = x+y Let Lets make table w? & P(W) P (W) W=P (W) 0.16 0.08 0.04 0.16 + 0.36 = 0.52 1.56 4.68 0.44 1.76 7.04 Total 3.4 11.E(X) - 1.6 var(x) = E(X²) - [E(X)]2 - 2.80 - 1.62 = 2.80 - 2.56. var(x) = 0.25 YPLY) , y2P(4) 0.20 0.20 0-20 1 Y PLY) 10-20 2

So variance of W = X + Y is same for both the methods.